OpenStudy (hlilly2413):

When you find the molarity of a titrant, do you add the initial and the final volume together and use that as the volume? (Example: A buret containing NaOH solution has an initial reading of 0.14 mL. 25.00 mL is 0.1000 M HCl is titrated to a phenolphthalein end point, which occurs at a final buret reading of 27.16mL NaOH. What is the molartiy of the NaOH titrant? I found the number of moles of NaOH, but am not sure what to use as the volume. Please help.

OpenStudy (hlilly2413):

My equation looks like this: \[NaOH (aq) + HCI (aq) \rightarrow H _{2}O (l) + NaCl (aq) \] and my work to find moles looks like this: \[25.00mL HCl \times \left( \frac{ 1LHCl }{ 1000mL HCl } \right)\times \left( \frac{ 0.1000MHCl }{ 1LHCl } \right)\] and then in continuation \[\times \left( \frac{ 1 mol NaOH }{ 1 mol HCl } \right) = 0.0025 moles NaOH \] Now, I know that in order to find the M of NaOH I need to do the However, I do not know which volume to use...if I add the 0.14 to the 27.16 or just use the 27.16 (obvi converted to L) . Any help is appreciated! Thanks!